OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x)=4/x g(x)=4/x

OpenStudy (anonymous):

\[f(g(x))=4/(4/x)\] 4/1 * x/4 cancel the 4s so you get x

OpenStudy (anonymous):

Since they're the same equation, you've already solved g((f(x)) and know that it's equal to x.

OpenStudy (anonymous):

So, x=1?

OpenStudy (anonymous):

No. x is equal to both f(g(x)) and g(f(x)) which can be considered in very simple terms the y values. The slope is 1 but x is not.

OpenStudy (anonymous):

What I gathered from your question is that you're only trying to prove that f(g(x)) is equal to g(f(x)) and that is done by solving each and comparing the results, both of which are x. Because they match, they're equal and according to your question that makes them inverses.

OpenStudy (anonymous):

So g(x) is the same as f(x) they both equal x?

OpenStudy (anonymous):

g(x) and f(x) are not equal to x. g(f(x)) and f(g(x)) are equal to x. Do you understand the difference between g(x) and g(f(x))?

OpenStudy (anonymous):

no :(

OpenStudy (anonymous):

Ok so if f(x) is the function of x and g(x) is a different function of x, f(g(x)) is the function of g(x). In other words, you're using the equation of g(x) to plug in for the x in the original f(x) equation. For example: f(x)=x+2 and g(x)=2x f(g(x))=(2x)+2

OpenStudy (anonymous):

So Both f(g(x)) and g(f(x)) are the same so that makes them inverses?

OpenStudy (anonymous):

By the definition you gave "Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x)=4/x g(x)=4/x" Yes. According to the question it can be proved that they're inverses if both f(g(x)) and g(f(x)) equal x, which you know they do because you solved both equations.

OpenStudy (anonymous):

Thank you so much, you are so smart. :)

OpenStudy (anonymous):

No problem! Not smart, just more experience so far. :) Let me know if you need help with anything else.

OpenStudy (anonymous):

Can you help me with a graph?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Just tag me in the question post.

OpenStudy (anonymous):

Determine if the function is one-to-one

OpenStudy (anonymous):

@christina18

OpenStudy (anonymous):

A function for which every element of the range of the function corresponds to exactly one element of the domain. In other words if for every x value there is only one y value. You can determine this by doing the vertical and horizontal line tests. Have you heard of them?

OpenStudy (anonymous):

That's where it crosses through two or more points on a graph with a horizontal/ vertical line?

OpenStudy (anonymous):

That's what they test, correct. Basically if you were to put your pencil flat across the graph horizontally and move it up across the graph and it never touches the graph in more than one place it passes. The vertical line test uses the same process except obviously vertical. To be one to one a graph must pass both tests. Does it?

OpenStudy (anonymous):

No, the horizontal line goes through more than one value

OpenStudy (anonymous):

Correct, so what does that mean?

OpenStudy (anonymous):

It's not a one to one :)

OpenStudy (anonymous):

Exactly! Nice job.

OpenStudy (anonymous):

Thank you so much :)

OpenStudy (anonymous):

No problem.